# Forecasting the future development in quality and value of professional football players for applications in team management

### v1

@@ 5.1 General population of players
The two loss functions show differences in performance as the linear models predict relatively worse with respect to the MAE than in terms of the RMSE compared to the other models. Consequently, the linear models have relatively few large errors, which corresponds to the players that are harder to predict. On the contrary, the linear models predicted worse for players that are easier to predict. This means that the linear models predicted well for the players whose development is less predictable, but relatively bad for the players that are relatively easy to predict.
Additionally, the kNN models, which have time-series information as features, predict better than the linear models. These results show that the time series of the player performance indicators contain important information for the development of the transfer value and that nonlinearities and interactions are involved.
5.2 Predictions per age
The RMSE values on the test set for each age are given in Figure 12. The loss generally decreases as players grow older. Because old players decrease in value, their values are smaller and more predictable, which explains the pattern. In general, no model is preferable based on the results of the young players.
Our models predict the development of the transfer value for young players with less accuracy, which is reflected by the high loss values for younger players. This indicates that the development of player value for young players is less predictable and this can be expected as the development in the player quality is found to be less predictable for younger players in the player quality prediction problem (section 4). Because the development of the transfer value depends on the development of a football player’s quality for young players, it can be assumed that the errors propagate in this model for player value. This explains why the errors for the young players are the largest.
@@ 5.2 Predictions per age
5.3 Prediction on important player groups
For the development of the transfer value, the four different important groups of players that are important for the application of the models are: high-quality players, high-value players, players with a large decrease in transfer value, and players with a large improvement in transfer value. The RMSE for these groups of players is visualized in Figure 14, Figure 14, Figure 16, and Figure 16. Similar to the prediction problem for player quality, the errors in the predictions are larger for these interesting groups of players.
@@ 5.3 Prediction on important player groups
The results indicate that the random forest model predicts best for the high-quality players, high-value players, and players with a large decrease. In contrast, the XGBoost model is better able to capture large increases in value than the random forest model. The performance of the random forest model on the players with a high player value is evidently better than the other models. Players often attain high transfer values for a short period in their careers when they perform well and are young. Consequently, players with high transfer values are expected to decrease in value after shortly obtaining high transfer values. The fact that the random forest model assigns large importance to the most recent transfer value suggests that the random forest can capture this pattern, which explains the better predictions on this group of players.
5.4 Feature importances
@@
The feature importances for the development of the player transfer value after min-max scaling are shown in Figure 17. The results indicate that the most important features for predicting player value development are the features describing the most recent transfer value and the developments within the last 6 and 12 months of the player transfer value. Additional time series information about the player quality (‘sciskill_diff_6m_ago’) also provides information about the development of the player transfer value. This implies that the time series information is most important in predicting future development, which is in line with the relatively good predictive performance of the kNN models that are based on time series information.
@@
It is also found that the month of the year is an important factor. This can be explained by the fact that the transfer value of a player developed differently at different times of the year. For example, a player’s transfer value tends to decrease when he only has 6 months of contract left because he could negotiate with other professional football clubs. As this commonly happens in the winter, the development of players’ transfer value is different depending on the time of the year.
@@ 6 Conclusion
This paper aims to find the most suitable supervised learning model for forecasting the development of player performance indicators one year ahead. Two prediction problems have been considered to study the predictive performance of linear, tree-based, and kNN-based models. The prediction problem of forecasting player quality (SciSkill) shows that the XGBoost model provides the most accurate predictions on the test set. This is especially the case for the general set of players, the players with the younger ages, and some interesting groups of players. Those interesting groups of players are the high-quality players and the players with either a large decrease or increase in football quality. The random forest model predicts second best with similar performances on the young players and most of several interesting groups of players. For the problem of predicting a football player’s development in quality, it is concluded that the best predictive performance is attained by the XGBoost model, followed closely by the random forest model.
The second prediction problem is about forecasting the development of the monetary values of football players (ETV) within the subsequent year. The results show that the most accurate predictions are obtained using the random forest model. This is mainly indicated by the loss values of the general test set and the groups of players with high quality, high value, or with a large increase or decrease in player value. Consequently, it is concluded that the random forest model provides the most accurate predictions of the development of a football player’s monetary value.
The aim was to assess the model quality not only based on the predictive performance of a model, but also to take explainability and uncertainty quantification into account in order to make the models more attractive for practitioners. As all trained models are explainable, the models now have to be assessed based on the availability of off-the-shelf methods for uncertainty quantification.
Linear regression has an underlying theory that gives prediction intervals. Although these prediction intervals are based on assumptions that did not hold for the prediction problems in this paper, such as the normality of errors, they do give an indicator for the uncertainty of the prediction. Similarly, the kNN models provide the neighbors, which is a group of similar data points. Uncertainty quantification can be obtained by taking the minimal and maximal values of the dependent values within these neighbors if the number of neighbors $k$ is of significant size. Lastly, the bagging procedure of the random forest model can be utilized to obtain uncertainty quantification for the predictions, as described by Wager et al. [47]. From these properties, it is concluded that the linear regression, random forest, and kNN-based models are favorable with respect to uncertainty quantification.
@@
In short, our results show that the predictive performance of the random forest has the second-best predictive accuracy for predicting the development of the player quality and that it provides the best predictive accuracy for the forecasting of the player value. Moreover, an off-the-shelf method for uncertainty quantification is available. Therefore, it is concluded that the random forest is the most suitable explainable machine learning model to predict the development of player performance indicators one year ahead.
7 Discussion
@@
By addressing the two prediction problems, two random forest models have been obtained that can predict the development of both the quality and the monetary value of football players. These models can be used to aid in transfer decisions. Suppose the manager of a football club has a long-term interest in a football player. If the model predicting the development of player quality indicates that a player will grow in quality, it means that a player is more interesting to a football club in the long term. If the prediction of the transfer value indicates an increase of the transfer value, it might be better to buy the player sooner rather than later. The two models can also be combined to give extra insights. Suppose a manager currently has a veteran player who is predicted to start decrease in quality, and he can buy a young player who is predicted to develop into a first-team player within the next year. Assume that the transfer value of the young player is is predicted to only slightly increase. In this case, it might be better to buy the young player one year later and then sell the veteran player, because the young player will only have a slightly higher transfer fee and the veteran player will have a better quality in the meantime. These examples illustrate the added value of our models to predict player development for the improvement of data-informed decision making.
@@
The models from this paper can also be used to complement existing methods in the literature. Pantuso and Hvattum [19] introduced a method to optimize transfer decisions based on indicators that describe a player’s quality and transfer value and their future values. Their methodology needs to know the ‘future’ values of the player quality and their transfer values. To solve this, they consider transfer situations from over a year prior so that the values of one year later were already known. Consequently, their model can only be applied to historical situations. By using our models to predict the future values in quality and transfer value, it is possible to obtain these predictions and optimize transfer decisions with a current-time application. This makes it possible to advise data-driven transfer decisions to optimize the squad in a real-life transfer period.
@@
An advantage of the explainable models in this research is that they provide methods to gain insight into important factors for prediction. The models for the prediction problems show that this can be done via feature importances similarly as done by Baouan et al. [39]. Because our models predict the difference between the current performance indicators and that of one year later, our research can also show the influences of the indicator itself. Our results show that the value of the indicator and the historical values of the indicators are the most important features, which adds to the knowledge obtained by Baouan et al. [39] that describes which features are important. Additionally, time-dependent variables like the period in the year and the months without games have been found to be important features in forecasting player development. This indicates that the time series of the indicators itself contains important information on the development of football players, although our findings also suggest that contextual information gives improved predictive performance of the models.
@@
In short, this paper studied explainable supervised models to predict player development via performance indicators. Two prediction problems were studied in which explainable models were trained to predict the development of both player quality and player value. It was found that the random forest model is the most suitable model for forecasting player development, because of the accurate predictions for both performance indicators combined with the method for uncertainty quantification arising from the bagging procedure.

### v2

@@ 4.2.1 General population of players
The two loss functions show differences in performance as the linear models predict relatively worse with respect to the MAE than in terms of the RMSE compared to the other models. It can be reasoned that the linear models have relatively few large errors, which corresponds to the players that are harder to predict. On the contrary, the linear models predicted worse for players who are easier to predict. This means that the linear models predicted well for the players whose development is less predictable, but relatively badly for the players who are relatively easy to predict.
Additionally, the kNN models, which have time-series information as features, predict better than the linear models. These results show that the time series of the player performance indicators contains important information for the development of the transfer value and that nonlinearities and interactions are involved.
4.2.2 Predictions per age
The RMSE values on the test set for each age are given in Figure 13. The results show that the loss values of the models are similar at all ages and that there is no model that outperforms the other models on players of a young age. Therefore, no model is favorable based on the prediction of young players.
@@ 4.2.2 Predictions per age
Moreover, the linear models perform worse on older players. This suggests that the patterns in the development of older players involve interaction effects or nonlinearities that were not captured by the linear models.
Generally, the models tend to predict the future development of player value worse for young players and better for older players. The development of a young player is less predictable, similarly as found for the prediction of player quality in subsection 4.1. This is due to the fact that the KPIs for young players are more volatile. On the other hand, the player values of older players are smaller and more predictable because old players tend to decrease in value. In this way, it can be explained why the models have large loss values for young players and small loss values for older players.
4.2.3 Prediction on important player groups
For the development of the transfer value, the four different important groups of players that have been identified as important for the application of the models are: high-quality players, high-value players, players with a large decrease in transfer value, and players with a large improvement in transfer value. The RMSE for these groups of players is visualized in Figure 15, Figure 15, Figure 17, and Figure 17. Similar to the prediction problem for player quality, the errors in the predictions are larger for these interesting groups of players.
@@ 4.2.3 Prediction on important player groups
The results in Figure 15 and Figure 17 show that the random forest model and kNN-based models perform best for the players of high quality and the players with a large decrease in value. The low loss values of the kNN-based models indicate that the time series features described in LABEL:Appendix:_Table_knn_features_player_value contain most of the important information to predict the development of these subsets of players. The good performance of the kNN also indicates that a local method is suitable for the prediction of these types of players. On the other hand, Figure 17 shows that, for players with a large increase, the kNN models perform relatively worse compared to the tree-based models. Similarly, the tree-based models also outperform the linear model on the subset of players with a large increase. These differences suggest that a large increase in player value is influenced by many different features and contains nonlinear patterns or interaction effects.
@@
Figure 15 shows that the performance of the random forest model on the players with a high player value is evidently better than that of the other models. Players often attain high transfer values for a short period in their careers when they show peak performance and are young. Consequently, players with high transfer values can sometimes be expected to decrease in value shortly after obtaining high transfer values. The random forest model seems to be the only model that captures this pattern, as will be indicated by the high feature importance of the current transfer value in subsubsection 4.2.4. This explains the significantly better performance of the random forest model on the high-value players.
Generally, the random forest model attains relatively low loss values on the interesting subgroups of players. This makes the random forest favorable over the other models. Additionally, the kNN-based models perform well on high-quality players and players with a large decrease, making them more suitable to predict star players on their decline.
4.2.4 Feature importances
@@
The feature importances for the development of the player transfer value after min-max scaling are shown in Figure 18. The results indicate that the most important features for predicting player value development are the features describing the most recent transfer value and the developments within the last 6 and 12 months of the player’s transfer value. Especially, the random forest model assigns a large value to the player value at the time of prediction (‘etv’), which explains its good performance on high-value players in Figure 15. Figure 18 also shows that additional time series information about the player quality (‘sciskill_diff_6m_ago’) provides information about the development of the player transfer value. This implies that the time series information is most important in predicting the future development, which is in line with the relatively good predictive performance of the kNN models that are based on time series information.
@@
Moreover, it is found that the month of the year is an important factor. This can be explained by the fact that the transfer value of a player develops differently at different times of the year. For example, a player’s transfer value tends to decrease when the player only has 6 months of contract left because he could negotiate with other professional football clubs. As this commonly happens in the winter and the Estimated Transfer Value model contains the indicator whether a player has less than 6 months on his contract, the development of players’ transfer value is different depending on the time of the year.
4.3 Predictive accuracy
The results showed that the random forest model attained the best loss values on the general population of players, the groups of players with high quality, high value, or with a large increase or decrease in player value. To conclude, it is found that the random forest model provides the most accurate predictions of the development of a football player’s monetary value.
@@ 5 Conclusion
This paper aims to find the most suitable supervised learning model for forecasting the development of player performance indicators one year ahead. Through the literature, it was found that linear regression, random forest, and kNN-based models are favorable based on their uncertainty quantification methods. Two prediction problems have been considered to study the predictive performance of linear, tree-based, and kNN-based models.
@@
The results show that the predictive performance of the random forest has the second-best predictive accuracy for predicting the development of the player quality and that it provides the best predictive accuracy for the forecasting of the player value. Moreover, an off-the-shelf method for uncertainty quantification is available. Therefore, it is concluded that the random forest is the most suitable explainable machine learning model to predict the development of player performance indicators one year ahead.
6 Implications and future directions
@@
By addressing the two prediction problems, two random forest models have been obtained that can predict the development of both the quality and the monetary value of football players. These random forest models can be used to aid in transfer decisions. Suppose the manager of a football club has a long-term interest in a football player. If the model predicting the development of player quality indicates that a player will grow in quality, it means that a player is more interesting to a football club in the long term. If the prediction of the transfer value indicates an increase in the transfer value, it might be better to buy the player sooner rather than later. The two models can also be combined to give extra insights. Suppose a manager currently has a veteran player who is predicted to start decreasing in quality, and he can buy a young player who is predicted to develop into a first-team player within the next year. Assume that the transfer value of the young player is predicted to only slightly increase. In this case, it might be better to buy the young player one year later and then sell the veteran player, because the young player will only have a slightly higher transfer fee, and the veteran player will have a better quality in the meantime. These examples illustrate the added value of our models to predict player development for the improvement of data-informed decision-making.
@@
The models from this paper can also be used to complement existing methods in the literature. Pantuso and Hvattum [19] introduced a method to optimize transfer decisions based on indicators that describe a player’s quality and transfer value and their future values. Their methodology needs to know the ‘future’ values of the player quality and their transfer values. To solve this, they consider transfer situations from over a year prior so that the values of one year later are already known. Consequently, their model can only be applied to historical situations. By using our models to predict the future values in quality and transfer value, it is possible to obtain these predictions and optimize transfer decisions with a current-time application. This makes it possible to advise data-driven transfer decisions to optimize the squad in a real-life transfer period.
@@
An advantage of the explainable models in this research is that they provide methods to gain insight into important factors for prediction. The models for the prediction problems show that this can be done via feature importance, similar as done by Baouan et al. [47]. Because our models predict the difference between the current performance indicators and those of one year later, our research can also show the influences of the indicator itself. Our results show that the value of the indicator and the historical values of the indicators are the most important features, which adds to the knowledge obtained by Baouan et al. [47] that describes which features are important. Additionally, time-dependent variables like the period in the year and the months without games have been found to be important features in forecasting player development. This indicates that the time series of the indicators themselves contains important information on the development of football players, although our findings also suggest that contextual information gives improved predictive performance of the models.
@@
In short, this paper studied explainable supervised models to predict player development via performance indicators. Two prediction problems were studied in which explainable models were trained to predict the development of both player quality and player value. It was found that the random forest model is the most suitable model for forecasting player development, because of the accurate predictions for both performance indicators, combined with the method for uncertainty quantification arising from the bagging procedure.